caro1717:

. A random variable r is normally distributed with a mean of 7 and a standard deviation of 1.5. Find the value of w so that P ( 8.8 < r < w) = .0800

jhonyy9:

@XioGonz any idea here ? pls.

XioGonz:

( 8.8< r<w ) = 0.08 ( r < w) -( r < 8.8) = 0.08 ( r<w) = 0.08 + (r < 8.8) (r<w) = 0.08 + ( z < (8.8-7)/1.5) = 0.08 + ( z < 1.2) = 0.08 + 0.88493 = 0.96493 (z < 1.811006) = 0.96493 But wait there's more..

XioGonz:

z = (w-m) 1.811006 = (w- 7)/1.5 w = 1.811006*1.5+7 So = 9.716509

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz ( 8.8< r<w ) = 0.08 ( r < w) -( r < 8.8) = 0.08 ok ty - but can you explain how you get this result of ... ? -(r<8,8) i m not sure that the student who asked the problem will can understanding it ... ( r<w) = 0.08 + (r < 8.8) (r<w) = 0.08 + ( z < (8.8-7)/1.5) = 0.08 + ( z < 1.2) = 0.08 + 0.88493 = 0.96493 (z < 1.811006) = 0.96493 But wait there's more.. \(\color{#0cbb34}{\text{End of Quote}}\)

XioGonz:

Find the value of w so that P ( 8.8 < r < w) = .0800 -from the question.